OpenStudy (anonymous):

hi i really need help with this calc question. find the extrema, points of inflection, intervals where incre/dec, intervals where concave up or down, and any vertical/horizontal asymp. y= 2x(x-3)^2

OpenStudy (danjs):

find y' and y'', then the calculus is over

OpenStudy (danjs):

you need to sketch the basic graph from this info?

OpenStudy (danjs):

notice it is a cubic function, it prolly turns around a couple times inc/dec

OpenStudy (danjs):

What you get for the first derivative? product and chain rule

OpenStudy (anonymous):

okay this is what i got

OpenStudy (anonymous):

as my first direvitive= 2(c-30^2 +2(x-3)x

OpenStudy (anonymous):

@DanJS also i need help with the second derivative.

OpenStudy (danjs):

man, sorry, i dont get notifications

OpenStudy (anonymous):

yes, me too

OpenStudy (anonymous):

okay so if my first part is correct which i do no tthink it is. plz tell me how to do the derivative

OpenStudy (danjs):

the function is a product of two terms, 2x and (x - 3)^2, recall the power rule [f*g] ' = first * derivative second + second * derivative first

OpenStudy (danjs):

and the chain rule for the (x-3)^2 term when you have to take it's derivative

OpenStudy (anonymous):

2(x-3)? so my first derivative is right?

OpenStudy (danjs):

\[y' = (2x)*2(x-3)^1*1 + (x-3)^2 * 2\] \[y' = 6*(x-3)*(x-1)\]

OpenStudy (anonymous):

and critical points there would be 3 and 1

OpenStudy (danjs):

cleans up good after expanding all that out and simplify

OpenStudy (danjs):

yes critical points to test where maybe the thing changes from increaing to decreasing, are where the slope is zero there (horizontal), y ' = 0, yep, when xis 3 or 1

OpenStudy (anonymous):

so what about the second derivitive?

OpenStudy (danjs):

x = 3, point y = 0 x = 1, y = 8 (3,0) and (1,8) you can add to the graph, possible change in directions

OpenStudy (danjs):



OpenStudy (danjs):

just pick any easy number to use inside each of those intervals between the critical points, and see if the derivative (slope) is positive or negative

OpenStudy (anonymous):

yes, that how i am doing it too. thank that so nice of you to help me. :) tysm. but i have a question.

OpenStudy (anonymous):

(2x)∗2(x−3)1∗1+(x−3)2∗2 how did u got 6(x-3)(x-1) from the equation above y′=6∗(x−3)∗(x−1)

OpenStudy (anonymous):

are u there?

OpenStudy (danjs):

sorry had to do somthing back

OpenStudy (danjs):

expand it all out, and combine like terms, then factor, i think , (i just used calculator)

OpenStudy (anonymous):

it's okay. i wiill do anything to get done with this problem today lol. okay so this is what i had 2((x-3)^2 +2(x+3)x)

OpenStudy (danjs):

how did you get that?

OpenStudy (anonymous):

wait let me attach a file

OpenStudy (danjs):

that equals 6(x^2+3),

OpenStudy (anonymous):

are u able to open it?

OpenStudy (danjs):

[2x] ' = 2 [(x-3)^2] ' = 2(x-3)^1 * 1

OpenStudy (anonymous):

did u open the link?

OpenStudy (danjs):

oh are you doin the second derivative

OpenStudy (anonymous):

no, i am talking about the first

OpenStudy (anonymous):

did u see if the first derivitive u got is right or not with the one i got from the website?

OpenStudy (danjs):

let me look at the site, pretty sure mine is right, it matches the graph

OpenStudy (anonymous):

yeah, i think u are right too but just making sure.

OpenStudy (danjs):

yeah that last answer is the same, just messy looking

OpenStudy (anonymous):

so yours is right

OpenStudy (danjs):

yeah, so is the website in an uglier form, you just copied it here wrong, so it wasnt good..

OpenStudy (anonymous):

but it's okay if i write yours?

OpenStudy (danjs):

yeah i would, that is probably what they want in your class or book

OpenStudy (anonymous):

okay. thanks! i still need help with the second derivitive

OpenStudy (danjs):

y = 2x(x-3)^2 y' = 6*(x-3)*(x-1) y'' = 12x - 24

OpenStudy (danjs):

using the chain rule, pretty simple, the derivative of those quantities is 1, and th equantity to the zero power is 1 or you can multiply it all out, and just use the power rule

OpenStudy (anonymous):

ao y "= 1

OpenStudy (anonymous):

so*

OpenStudy (danjs):

y'' = 6 * [ (x-3) * 1*(x-1)^0 *1 + (x-1)* 1 *(x-3)^0 *1 ]

OpenStudy (danjs):

that is the chain rule, more typing than thinking , 6*[(x-3) + (x-1)] 12x - 24

OpenStudy (danjs):

If you expand y' first, you get y ' = 6x^2 - 24x + 18 y ' ' = 12x - 24 (easy power rule in that case)

OpenStudy (anonymous):

thanks but the answer whould be 12(x-2) critical point at 2

OpenStudy (danjs):

lol yeah, factor a 12 out

OpenStudy (anonymous):

for the y"

OpenStudy (anonymous):